# Realistic Cost to Execute Practical Quantum Circuits using Direct Clifford+T Lattice Surgery Compilation

### v2

@@ 4.3. Total Resource Costs and Calculation of Error Rates
Using this information, we may estimate error rates for the logical, distillation, and storage parts of the computation. Given a code distance $d$ and a hardware-dependent physical error rate $p$, we calculate the logical error rate of a surface code patch as $P(d)=0.03(\frac{p}{0.01})^{\frac{d+1}{2}}$, where the pre-factor and the threshold value are taken from Ref. (Fowler and Gidney, 2018). We consider two different sets of estimates for hardware gate infidelities that can be selected through the param_type option in the resource analysis script (for --param_type current, $p_{1}=4\times 10^{-5}$ and $p_{2}=p_{I}=3\times 10^{-3}$ (Quantinuum, 2023); for --param_type projected, we modify these using recent demonstrations: $p_{I}=1\times 10^{-4}$ (An et al., 2022) and $p_{2}=6\times 10^{-4}$ (Clark et al., 2021)) [^22] These numbers are based upon current and projected estimates for the infidelity of single-qubit gates ($p_{1}$), two-qubit gates ($p_{2}$), and state initialization ($p_{I}$) in trapped-ion hardware specifically; one could easily modify the script to include parameters pertaining to other hardware modalities. The resource estimates reported in Sec. 5 use the projected parameter set.. Since syndrome extraction circuits are primarily made of CNOT gates (Fowler et al., 2012), we use $p=p_{2}$ in the formula for $P(d)$. Further, since our analysis requires injected magic states as input to the first level of distillation, we calculate the injected state error according to the formula from Ref. (Lao and Criger, 2022) (this is a refinement of the state injection protocol of Ref. (Li, 2015) for the rotated surface code), which is $P_{I}=\frac{3}{5}p_{2}+p_{I}+\frac{2}{3}p_{1}$.
@@ 5.1. Logical and Fault-Tolerant Resource Estimates
In this section, we present resource estimates for large practical quantum circuits, namely those corresponding with one Trotter step of an efficient circuit implementation of the Trotter-Suzuki ground-state estimation (GSE) algorithm that was proposed in Ref. (Kornell and Selinger, 2023). To compile this algorithm into the gate set detailed in Sec. 2.2, we have used Quipper (Green et al., 2013) together with with LinguaQuanta [^24] LinguaQuanta is a Quipper to QASM converter that can be found at https://github.com/onestruggler/qasm-quipper.. From these circuits, we compile lattice surgery circuits and estimate resources for 5, 7, 9, 11, and 13 orbitals. See Table 2 for details on the numbers of logical qubits, logical gates, sliced LLI compiled to by the LSC [^25] The number of sliced LLI differs from the number of LLI in that where an LLI appears on multiple slices (as is the case in BellBasedCNOT and RotateSingleCellPatch) it is multiply counted. This number does not include local instructions., and MagicStateRequest LLI (equivalent to the number of T/$\mathrm{T}^{\dagger}$ gates in the circuit) for each of these circuits, as well as the LSC compile time. The largest circuit that we compile has around ten million gates. We note that the bottleneck to our pipeline is in algorithm compilation to our input gate set rather than in lattice surgery compilation, which in the 13 orbital case took a little under nine hours. Our resource analysis script, the functionality of which was detailed in Sec. 4, uses one of three resource optimization approaches (‘default’, ‘add-warms’, or ‘min-storage’) to find the minimum value for an input cost function. We present results using all three of these approaches to optimize space-time volume, for which we use $n_{total}\times\tau_{total}\times d_{2}^{3}$ as a proxy. Our main results can be found in Tables 34, and 5, which show optimized parameter values ($d_{1}$, $d_{2}$, $N$, and $w_{total}$) as well as resource estimates: $n_{total}$ (in tiles), $\tau_{total}$ (in slices), $n_{total}\times\tau_{total}$ (in tile-slices), $V_{total}$ (in tile-slices), $n_{total}\times\tau_{total}\times d_{2}^{3}$ (our proxy for space-time), and $\epsilon_{total}$ (total error rate).
@@
Orbitals Logical Qubits Gates Sliced LLI MagicStateRequest LLI LSC Compile Time (s)
5 23 2.11e+05 2.70e+06 8.31e+04 2.57e+02
7 31 7.90e+05 1.00e+07 3.10e+05 1.63e+03
9 39 2.17e+06 2.74e+07 8.47e+05 7.29e+03
11 47 4.88e+06 6.16e+07 1.91e+06 2.15e+04
13 55 9.62e+06 1.21e+08 3.75e+06 3.11e+04
@@
5 7 17 6 1804 1.15e+04 2.75e+05 3.16e+09 1.61e+09 1.55e+13 1.55e-03
7 15 19 9 2004 4.20e+04 8.32e+05 3.49e+10 1.72e+10 2.40e+14 9.72e-04
9 17 19 11 2604 8.01e+04 1.93e+06 1.54e+11 7.93e+10 1.06e+15 4.54e-03
11 11 19 11 5404 6.16e+04 3.75e+06 2.31e+11 9.56e+10 1.58e+15 5.59e-03
13 15 19 14 2404 5.01e+04 6.47e+06 3.24e+11 1.26e+11 2.22e+15 6.79e-03
Table 5. Resource estimates and optimal parameter values for our min-storage resource estimation approach (factories are switched on to match consumption rate.)
@@
First, we note that while we used $\epsilon_{thr}=0.01$ (representing a success rate of 99%) as a constraint to our optimization protocol, our results typically do not saturate this bound. Instead, as the number of orbitals (and thus the size of the circuit) increases and the code distances remain constant, the total error rate increases until the code distance needs to increment due to the total error passing the threshold, at which point $\epsilon_{total}$ drops. Second, we notice that while $\tau_{total}$ values did not substantially change between resource optimization approaches (since the primary contribution to this quantity is the number of slices from the LSC output, which is not changing), the $n_{total}$ values change dramatically. The most important result from Tables 34, and 5 is the scaling of space-time volume with respect to the number of orbitals, which is highlighted in Fig 7. The superiority of the min-storage approach over the default and add-warms approaches for optimizing this quantity is evident. Since the primary reason that space-time is changing between approaches is that $n_{total}$ is changing, we now investigate how much of the quantum computer is being dedicated to each task (logical computation, magic state distillation, or magic state storage) between the three resource optimization approaches. As anticipated in Sec. 4, we found that most of the spatial resources required by the computation in the default and add-warms approaches come from magic state storage. To see this most vividly, Fig. 8(a) and (b) show that storage requires $\sim 99\%$ of the hardware tiles in the former case and $\sim 90\%$ in the latter, while it contributes the least to the overall spatial cost in the min-storage approach [Fig. 8(c)]. In this latter case, magic state distillation becomes the dominant cost, bringing us into the regime that people typically consider when they neglect magic state storage. As mentioned in Sec. 4, the min-storage approach involves allocating more space to distillation factories than is typically utilized within a given distillation cycle (the min-storage approach requires more than four times as many factories as the add-warms approach for 13 orbitals); therefore, our results show that minimizing resources dedicated to magic state distillation is not always beneficial, and that a simplifying assumption of a constant rate of magic state consumption to determine the number of factories (which bears the most similarity to our add-warms approach) can be truly disastrous due to the cost of storing excess states during lower-consumption sections of the circuit.
@@
Figure 9. Profile of total requested magic states (a) and number of reserved magic states at the end of each distillation cycle for three resource estimation approaches (b)-(d) for the 13-orbital ground state estimation circuit.
To make this last statement more concrete, Fig. 9(a) shows the cadence of magic state requests (coarse-grained by distillation cycle) for the 13-orbital GSE circuit. Looking at the profile of magic state requests as a function of distillation cycle, we see that the rate of requests is slightly higher in the beginning of the circuit than in the bulk of the circuit; other irregularities in the rate are masked by the large number of requests being made. For comparison, Figs. 9(b)-(d) show the number of magic states in reserve as a function of distillation cycle for each of the three resource optimization approaches. Because our methodology ensures that, in all three approaches, magic state supply keeps up with demand at all times, the need for a higher production rate at the beginning can result in a growing excess of magic states during the bulk of the circuit until production halts once every needed magic state has been produced; that is, if the number of factories that are switched ‘on’ is constant. This is especially apparent for the default approach in Fig. 9(b), but is also apparent for the add-warms approach in Fig. 9(c). In the latter case, adding warm-up cycles so that a larger bank of magic states exists at the beginning of the circuit does mitigate this problem to a great extent (the maximum number of states in reserve is about twenty times lower in this case), but it can not overcome the problem of having large numbers of excess magic states at various points in the circuit due to irregularities in consumption rate. Conversely, it can be seen for the min-storage approach in Fig. 9(d) that the number of states in reserve during any given distillation cycle is minimal, as expected by construction of that approach.
@@ 5.2. Trapped-Ion Resource Estimates
The resource estimates presented in the former section are in terms of hardware tiles and logical time-slices, which are abstract quantities from the fault-tolerant compilation layer. As of yet, the only information from the hardware layer that has entered into our calculation is the physical error rates (we have used the projected parameter set noted in Sec. 4). To obtain resource estimates in real terms (space and time) we will consider the cost to execute the same GSE circuit on a trapped-ion processor. To do this, we must consider the time to execute the slowest instruction from our local instruction set, as this will define the execution time for one slice, as well as the spatial requirement for a single hardware tile. To obtain these quantities, we make use of the Trapped-Ion Surface Code Compiler (TISCC) from Ref. (LeBlond et al., 2023a), which generates time-resolved circuits for the same set of local lattice surgery instructions considered in this paper. We will not consider details of the trapped-ion hardware model employed by TISCC here, and opt to simply state the final results. Using trapped-ion gate durations derived from the present literature, TISCC estimates a single round of error correction to require $9.613\times 10^{-3}$ s. Our best resource estimates for the 13-orbital GSE circuit are from Table 5, so we will use $d_{2}=17$, which yields the time cost of a logical time-slice to be $\tau_{0}=1.63\times 10^{-1}s$ [^26] We did not consider decoding time to dominate the cost of a logical time-step because the rate of syndrome extraction for trapped-ion systems appears to be slower than state-of-the art decoding algorithms (see benchmarking for PyMatching in Ref. (Delfosse et al., 2023)). . Querying TISCC further to estimate the spatial cost of a hardware tile at $d=17$, we find $n_{0}=9.15\times 10^{-4}m^{2}$. Combining these numbers with the results from Table 5, we find that the total time to run the circuit is $t=\tau_{total}\times\tau_{0}=1.05\times 10^{6}s$, which is just over twelve days, and that the total area needed to support the hardware grid is $A=n_{total}\times n_{0}=1.93m^{2}$.
@@ 6. Conclusion and Outlook
In this work, we have presented a resource estimation pipeline for practical quantum circuits that relies on explicit lattice surgery compilation within the fault-tolerant layer to obtain an accurate number of logical time-steps and active volume as well as a cadence of fulfilled magic state requests by time-step, all of which significantly improve the our ability to estimate resources relative to existing approaches. In order to implement our methodology, we have heavily revised the Lattice Surgery Compiler (Watkins et al., 2023) to include modified instruction sets, a local compilation layer, improved layout generation, and DAG-based parallelism, among other things. We used our pipeline to estimate resources for ground-state estimation circuits up to 13 orbitals within three different resource optimization approaches that explored trade-offs related to magic state distillation and storage, showing that assuming a constant (yet minimal, under the condition that there are always at least enough magic states in storage to fulfill all requests during a given distillation cycle) number of active magic state factories causes an unreasonable amount of space dedicated to storage in the quantum computer. Lastly, we showed how to combine our fault-tolerant resource estimates with results from a surface code compiler for trapped ion processors (TISCC) (LeBlond et al., 2023a) to obtain resource estimates in terms of space and time.
The work presented here relies on an abstract model of quantum computing hardware and is therefore agnostic to modality so long as the local, tile-based instruction set that has been used is desirable, though we understand that other instruction sets may be preferable for different modalities (see, for example, a surface code instruction set specific to neutral atom systems (Viszlai et al., 2023)). While in combination with Ref. (LeBlond et al., 2023a), the work presented here aids in the development of an end-to-end compilation and resource estimation pipeline for trapped-ion processors in which the final result has been verified to be correct using simulations of those systems, our pipeline was designed to be extensible to other approaches.
@@
At this stage of development in fault-tolerant compilation and resource estimation, we do not expect any resource estimates to be accurate to what a future system could achieve, as numbers can vary wildly according to different approaches, even just in considering trade-offs between magic state distillation and storage as we have done here. Our aim was instead to introduce a resource estimation pipeline with improved realism over other approaches through explicit lattice surgery compilation that accounts for circuit-level parallelism and the management of magic states. We anticipate that future work could better characterize the trade-space by further expanding the functionality of the LSC to include different compilation strategies and/or by considering integration with different hardware compilers like TISCC.

### v4

@@ 4.3. Total Resource Costs and Calculation of Error Rates
Using this information, we may estimate error rates for the logical, distillation, and storage parts of the computation. Given a code distance $d$ and a hardware-dependent physical error rate $p$, we calculate the logical error rate of a surface code patch per logical time-slice as $P(d)=0.1d(\frac{p}{0.01})^{\frac{d+1}{2}}$, where the pre-factor and the threshold value are taken from Ref. (Fowler and Gidney, 2018). We consider two different sets of estimates for hardware gate infidelities that can be selected through the param_type option in the resource analysis script (for --param_type current, $p_{1}=4\times 10^{-5}$ and $p_{2}=p_{I}=3\times 10^{-3}$ (Quantinuum, 2023); for --param_type projected, we modify these using recent demonstrations: $p_{I}=1\times 10^{-4}$ (An et al., 2022) and $p_{2}=6\times 10^{-4}$ (Clark et al., 2021)) [^23] These numbers are based upon current and projected estimates for the infidelity of single-qubit gates ($p_{1}$), two-qubit gates ($p_{2}$), and state initialization ($p_{I}$) in trapped-ion hardware specifically; one could easily modify the script to include parameters pertaining to other hardware modalities. The resource estimates reported in Sec. 5 use the projected parameter set.. Since syndrome extraction circuits are primarily made of CNOT gates (Fowler et al., 2012), we use $p=p_{2}$ in the formula for $P(d)$. Further, since our analysis requires injected magic states as input to the first level of distillation, we calculate the injected state error according to the formula from Ref. (Lao and Criger, 2022) (this is a refinement of the state injection protocol of Ref. (Li, 2015) for the rotated surface code), which is $P_{I}=\frac{3}{5}p_{2}+p_{I}+\frac{2}{3}p_{1}$.
@@ 5.1. Logical and Fault-Tolerant Resource Estimates
In this section, we present resource estimates for large practical quantum circuits, namely those corresponding with one Trotter step of an efficient circuit implementation of the Trotter-Suzuki ground-state estimation (GSE) algorithm that was proposed in Ref. (Kornell and Selinger, 2023). To compile this algorithm into the gate set detailed in Sec. 2.2, we have used the Quipper programming language (Green et al., 2013) together with LinguaQuanta (Wesley, 2024), a Quipper-to-QASM transpiler. We use Quipper’s default precision of $10^{-30}$ when decomposing rotations into Clifford+T gates. From these circuits, we compile lattice surgery circuits and estimate resources for 5, 7, 9, 11, and 13 orbitals. See Table 3 for details on the numbers of logical qubits, logical gates, logical T/$\mathrm{T}^{\dagger}$ gates (equivalent to the number of MagicStateRequest LLI), compiled slices, sliced LLI processed by the LSC [^25] The number of sliced LLI differs from the number of LLI in that where an LLI appears on multiple slices (as is the case in BellBasedCNOT and RotateSingleCellPatch) it is multiply counted. This number does not include local instructions., and the LSC compile time. The largest circuit that we compile has around ten million gates and requires a little under fifteen hours to compile [^26] While the compilation tasks were performed on relatively old (Skylake) CPUs available through the CADES cluster at ORNL, these compilation times highlight the need for advances in logical circuit decomposition to make practical quantum computing scalable, as mentioned in Sec. 3.1.. Our resource analysis script, the functionality of which was detailed in Sec. 4, uses one of three resource optimization approaches (‘default’, ‘add-warms’, or ‘min-storage’) to find the minimum value for an input cost function. We present results using all three of these approaches to optimize space-time volume, for which we use $n_{total}\times\tau_{total}\times d_{2}^{3}$ as a proxy. Our main results can be found in Tables 45, and 6, which show optimized parameter values ($d_{1}$, $d_{2}$, $N$, and $w_{total}$) as well as resource estimates: $n_{total}$ (in tiles), $\tau_{total}$ (in slices), $n_{total}\times\tau_{total}$ (in tile-slices), $V_{total}$ (in tile-slices), $n_{total}\times\tau_{total}\times d_{2}^{3}$ (our proxy for space-time), and $\epsilon_{total}$ (total error rate).
@@
Orbitals Logical Qubits Gates T/$\mathbf{T}^{\dagger}$ Gates Slices Sliced LLI LSC Compile Time (s)
5 23 2.11e+05 8.31e+04 2.31e+05 2.70e+06 5.97e+02
7 31 7.90e+05 3.10e+05 7.45e+05 1.00e+07 2.84e+03
9 39 2.17e+06 8.47e+05 1.76e+06 2.74e+07 1.02e+04
11 47 4.88e+06 1.91e+06 3.39e+06 6.16e+07 2.14e+04
13 55 9.62e+06 3.75e+06 5.70e+06 1.21e+08 5.26e+04
@@
5 7 19 6 1786 1.13e+04 2.63e+05 2.97e+09 1.51e+09 2.03e+13 1.66e-03
7 7 21 7 2931 4.07e+04 7.98e+05 3.25e+10 1.57e+10 3.01e+14 1.18e-03
9 21 21 12 2864 8.71e+04 1.84e+06 1.60e+11 8.40e+10 1.48e+15 5.96e-03
11 15 21 13 3942 6.18e+04 3.48e+06 2.15e+11 8.28e+10 1.99e+15 5.91e-03
13 13 23 14 2314 9.80e+04 5.75e+06 5.63e+11 3.01e+11 6.85e+15 1.48e-03
Table 6. Resource estimates and optimal parameter values for our min-storage resource estimation approach (factories are switched on to match consumption rate).
@@
First, we note that while we used $\epsilon_{thr}=0.01$ (representing a success rate of 99%) as a constraint to our optimization protocol, our results typically do not saturate this bound. Instead, as the number of orbitals (and thus the size of the circuit) increases and the code distances remain constant, the total error rate typically increases until the code distance needs to increment due to the total error passing $\epsilon_{thr}$, at which point $\epsilon_{total}$ drops [^27] Deviations from this trend in our results seem to stem from the complexity that the two-level factory design adds to the trade space.. Second, we notice that while $\tau_{total}$ values did not substantially change between resource optimization approaches (since the primary contribution to this quantity is the number of slices from the LSC output, which is not changing), the $n_{total}$ values change dramatically. This influences the most important result from Tables 45, and 6, which is the scaling of space-time volume with respect to the number of orbitals, shown in Fig 8. The superiority of the min-storage approach over the default and add-warms approaches for optimizing space-time volume is evident. Since the primary reason that space-time is changing between approaches is that $n_{total}$ is changing, we now investigate how much of the quantum computer is being dedicated to each task (logical computation, magic state distillation, or magic state storage) between the three resource optimization approaches. As anticipated in Sec. 4, we found that most of the spatial resources required by the computation in the default and add-warms approaches come from magic state storage. To see this most vividly, Fig. 9(a) and (b) show that storage requires $\sim 99.8\%$ of the hardware tiles in the former case and $\sim 97.4\%$ in the latter, while it contributes the least to the overall spatial cost in the min-storage approach [Fig. 9(c)]. In this latter case, magic state distillation becomes the dominant cost, bringing us into the regime that people typically consider when they neglect magic state storage. However, as mentioned in Sec. 4, the min-storage approach involves allocating more space to distillation factories than is typically utilized within a given distillation cycle (the min-storage approach requires three times as many factories as the add-warms approach for 13 orbitals); therefore, our results show that minimizing resources dedicated to magic state distillation is not always beneficial, and that doing so while holding magic state production constant can be disastrous in real circuits with variable consumption rates due to the cost of storing excess states produced during lower-consumption sections of the circuit.
@@
Figure 10. Profile of total requested magic states coarse-grained by distillation cycle (a) and number of reserved magic states at the end of every fiftieth distillation cycle for three resource optimization approaches (b)-(d) for the 13-orbital ground state estimation circuit. Since the number of slices per distillation cycle differs by approach, the profile from the default approach is used in (a), though the profile is qualitatively similar in the other cases.
To make this last statement more concrete, Fig. 10(a) shows the cadence of magic state requests (coarse-grained by distillation cycle) for the 13-orbital GSE circuit in the default approach. Looking at the profile of magic state requests as a function of distillation cycle, we see that the rate of requests is slightly higher in the beginning of the circuit than in the bulk of the circuit; other irregularities in the rate are masked by the large number of requests being made. For comparison, Figs. 10(b)-(d) show the number of magic states in reserve as a function of distillation cycle for each of the three resource optimization approaches. Because our methodology ensures that, in all three approaches, magic state supply keeps up with demand at all times, the need for a higher production rate at the beginning can result in a growing excess of magic states during the bulk of the circuit until production halts once every needed magic state has been produced; that is, if the number of factories that are switched ‘on’ is constant. This is especially apparent for the default approach in Fig. 10(b), but is also apparent for the add-warms approach in Fig. 10(c). In the latter case, adding warm-up cycles so that a larger bank of magic states exists at the beginning of the circuit does mitigate this problem to a great extent (the maximum number of states in reserve is about twenty times lower in this case), but it can not overcome the problem of having large numbers of excess magic states at various points in the circuit due to irregularities in consumption rate. In this case, a large bank produced at the beginning of the circuit, once depleted by the high-consumption portion of the circuit, gets re-built during lower-consumption portions until every needed magic state has been produced. Conversely, it can be seen for the min-storage approach in Fig. 10(d) that the number of states in reserve during any given distillation cycle is minimal, as expected by construction of that approach.
@@ 5.2. Trapped-Ion Resource Estimates
The resource estimates presented in the former section are in terms of hardware tiles and logical time-slices, which are abstract quantities from the fault-tolerant compilation layer. As of yet, our method has been entirely hardware agnostic except for the input of physical error rates (we have used the projected parameter set noted in Sec. 4) [^28] While it is expected that fault-tolerance schemes will be co-designed with hardware architectures, we have opted for a general-purpose scheme.. To obtain resource estimates in real terms (space and time) we will consider the cost to execute the same GSE circuit on a trapped-ion processor. To do this, we must consider the time to execute the slowest instruction from our local instruction set, as this will define the execution time for one slice, as well as the spatial requirement for a single hardware tile. To obtain these quantities, we make use of the Trapped-Ion Surface Code Compiler (TISCC) from Ref. (LeBlond et al., 2023a), which generates time-resolved circuits for the same set of local lattice surgery instructions considered in this paper. We will not consider details of the trapped-ion hardware model employed by TISCC here, and opt to simply state the final results. Using trapped-ion gate durations derived from the present literature, TISCC estimates a single round of error correction to require $9.613\times 10^{-3}$ s. Our best resource estimates for the 13-orbital GSE circuit are from Table 6, so we will use $d_{2}=19$, which yields the time cost of a logical time-slice to be $\tau_{0}=1.83\times 10^{-1}s$ [^29] We did not consider decoding time to dominate the cost of a logical time-step because the rate of syndrome extraction for trapped-ion systems appears to be slower than state-of-the art decoding algorithms (see benchmarking for PyMatching in Ref. (Delfosse et al., 2023)). . Querying TISCC further to estimate the spatial cost of a hardware tile at $d=19$, we find $n_{0}=1.13\times 10^{-3}m^{2}$. Combining these numbers with the results from Table 6, we find that the total time to run the circuit is $t=\tau_{total}\times\tau_{0}=1.04\times 10^{6}s$, which is just over twelve days, and that the total area needed to support the hardware grid is $A=n_{total}\times n_{0}=3.71m^{2}$.
@@ 6. Conclusion and Outlook
In this work, we have presented a resource estimation pipeline for practical quantum circuits that relies on explicit lattice surgery compilation to obtain an accurate number of logical time-steps, an accurate active volume, and a profile of fulfilled magic state requests by time-step, all of which significantly improve the realism of our estimates relative to existing approaches. In order to implement our methodology, we have heavily revised the Lattice Surgery Compiler (Watkins et al., 2024) to include modified instruction sets, a local compilation layer, improved layout generation, and DAG-based parallelism, among other things. We utilized three different resource optimization schemes within our pipeline to estimate resources for ground-state estimation circuits up to 13 orbitals to explore trade-offs related to magic state distillation and storage. We showed that assuming a constant (yet minimal, under the condition that there are always at least enough magic states in storage to fulfill all requests during a given distillation cycle) number of active magic state factories causes an unreasonable amount of space dedicated to storage in the quantum computer. Lastly, we showed how to combine our fault-tolerant resource estimates with results from a surface code compiler for trapped ion processors (TISCC) (LeBlond et al., 2023a) to obtain resource estimates in terms of space and time.
The work presented here relies on an abstract model of a quantum computer and is therefore hardware-agnostic so long as the local, tile-based surface code instruction set that has been used is desirable, though we understand that other instruction sets may be preferable for some hardware architectures (see, for example, a surface code instruction set specific to neutral atom systems (Viszlai et al., 2023)). While in combination with Ref. (LeBlond et al., 2023a), the work presented here aids in the development of an end-to-end compilation and resource estimation pipeline for trapped-ion processors in which the final result has been verified to be correct using simulations of those systems, our pipeline was designed to be extensible to other approaches.
@@
At this stage of development in fault-tolerant compilation and resource estimation, we do not expect any resource estimates to perfectly predict what a future system could achieve, as numbers can vary wildly according to different approaches; see for example, the trade-offs between magic state distillation and storage that we have considered in this work. Nonetheless, we believe our resource estimation pipeline achieves improved realism over other approaches through explicit lattice surgery compilation that accounts for circuit-level parallelism and the management of magic states. We anticipate that future work could better characterize the trade-space by further expanding the functionality of the LSC to include different compilation strategies and/or by considering integration with different hardware compilers like TISCC.